OpenStudy (anonymous):

Mario`s company makes unusually shaped imitation gemstones. One gemstone had 12 faces and 10 vertices. How many edges did the gemstone have? A) 23 B) 22 C) 25 D) 20

OpenStudy (mathstudent55):

Believe it or not, there is a formula for this. It's called Euler's Polyhedral Formula. For E = edges, F = faces, and V = vertices, the formula is V - E + F = 2

OpenStudy (anonymous):

your joking right?

OpenStudy (anonymous):

then, it must be option D Edges = 20, because 10 - E + 12 =2 => 22 - E = 2 => 22 - 2 + E => 20 + E

OpenStudy (anonymous):

Because you can`t solve for that because I don not know E

OpenStudy (mathstudent55):

F = 12, V = 10 V - E + F = 2 10 - E + 12 = 2 -E + 22 = 2 -E = -20 E = 20

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

haha, we had F = 12, and V = 10 :)

OpenStudy (anonymous):

haha right (: I was trying to subtract both sides instead of adding the 10 and 12...silly me!

OpenStudy (anonymous):

@mathstudent55 is a big help! I greatly appreciate it!

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

But, u got it :), yeah he is